# Search for a diffuse flux of photons with energies above tens of PeV at the Pierre Auger Observatory

### v2

@@ 1 Introduction
In this study, we extend the search for primary photons done at the Auger Observatory towards lower energies, reaching down to $5\times 10^{16}$ eV. The extension is made possible by a dense array of $19$ WCDs deployed over 2 km2 and the direct measurements of the air-shower muonic component by an array of buried scintillators deployed in the same region, as discussed in Section 2. The Monte Carlo-driven reconstruction of photon-initiated events and the definition of a unique energy scale to treat both photon and hadronic events are presented in Section 3. The photon-hadron discrimination is performed using an observable, $M_{b}$, defined as the event-wise weighted sum of the muon density measurements. The power of this photon-hadron discriminator is presented in-depth in Section 4. The selection criteria considered in data are described in Section 5. The results of the photon search are presented in Section 6.
@@ 2 Detection systems and data
In the configuration used in this work, eleven stations of the SD-433 array are co-located with stations of the Underground Muon Detector (UMD), as illustrated in Fig. 1, left. Each UMD station consists of three modules, each containing $64$ plastic scintillator bars measuring 400 cm in length, 4 cm in width and 1 cm in thickness providing a sensitive area of $\sim$10 m2 [42, 43]. The central station on the array’s western edge contains six modules, providing a sensitive area of $\sim$50 m2, compared to the standard 30 m2 area in all other stations. All modules are buried under 2.3 m of soil equivalent to a vertical shielding of 540 g/cm2, that filters out most of the particles produced in the air except muons with kinetic energies $\gtrsim$1 GeV. When a charged particle passes through the module, the scintillation material emits photons that are collected by a 1.2 mm wavelength-shifter optical fiber and conducted to an array of $64$ silicon photomultipliers (SiPMs) Hamamatsu S13081-050CS [44]. Front-end electronics convert the analog pulses of each SiPM into binary signals, effectively translating the current pulses into a sequence of digital boolean samples. These output signals are sampled by a Field-Programmable Gate Array at 320 MHz, corresponding to a sampling interval of 3.125 ns, resulting in a binary trace of $2048$ bits stored in the front-end memory. The back-end electronics handles all calibration, control, and monitoring tasks [45]. Additionally, the surface electronics, common to all UMD station modules, interfaces with the SD electronics to check for a trigger and retrieve and transfer the binary traces upon an event data request.
Figure 1: Left: A schematic view of the SD-433 array. Solid black positions are equipped with buried scintillators. Events acquired by the three highlighted hexagons are considered for the photon search. Right: One of the events acquired on February 27, 2021, part of the selected data set. The reconstructed energy is $(4.1\pm 0.1)\times 10^{17}$ eV, obtained with a dedicated data energy scale [38], with a zenith angle of $(26.3\pm 0.3)$. Dashed lines indicate the fitted lateral distribution functions (see Section 3).
@@
To reconstruct the incoming direction of the primary particle and the lateral distribution function (LDF) [48] of the air showers we use the signals from the WCDs. The LDF is described by the Nishimura-Kamata-Greisen (NKG) function [49, 50] from which we extract the shower size, $S(r_{\text{ref}})$, defined as the value at an optimal distance, $r_{\text{ref}}$. For an array spacing of 433 m, $r_{\text{ref}}$ is chosen as 250 m to minimize the uncertainties of the signal due to the imperfect knowledge of the functional form of the LDF in individual events [51]. The shower size, $S(250)$, corrected for the attenuation in the atmosphere serves as an observable that can be calibrated against the primary energy. Although for most events detected with SD-433 (hadronic showers) a data-driven energy calibration is employed [38, 52], the energy calibration for photon events is instead performed using detailed Monte Carlo simulations, as discussed in Section 3.
@@
The photon search strategy designed for this work relies on the SD-433 data to measure the air showers, providing information on the primary energy and shower geometry, and on the direct measurement of the air-shower muon content, as recorded by the UMD stations. To ensure a sufficient sampling of the air-shower muon content, leading to a suitable discrimination power as discussed in Section 5, we select events acquired by the three highlighted hexagons of Fig. 1, left. The data for this analysis were recorded from December 17, 2020, when the central UMD stations in the three employed hexagons were commissioned, to March 31, 2022, before the upgrade of the $19$ stations with new electronics [53].
@@
The flux of events acquired by the SD hexagons is expected to remain constant, apart from a negligible seasonal modulation. As the number of detected events follows Poisson statistics, identifying unstable periods in the data acquisition of the SD-433 involves searching for time intervals between consecutive events incompatible with the Poisson expectation [^1] For example, consecutive events separated more than 1.1 hours when the seven hexagons of the array are operative are not compatible with a Poisson probability at a $99\%$ confidence level and hence identified as an unstable period.. The same methodology is employed for the UMD modules, based on the rate of 6T5 events acquired by the SD-433 during which a UMD module is registered as active. Unstable periods for each UMD module are identified by analyzing the time intervals between consecutive appearances in the data. After subtracting unstable periods for both detection systems based on the arrival time of 6T5 events, the final data set is composed of 5T5 and 6T5 events recorded during ${\sim}15.5\,\mathrm{months}$, equivalent to $8\,\mathrm{months}$ of data acquisition with the three hexagons operating under ideal conditions. The final data set is composed of 15,919 events with energies above $10^{16.7}$ eV and zenith angles up to $52\,\mathrm{{}^{\circ}}$. The energy threshold is derived with the Monte Carlo study presented in Section 3.4, while the zenith angle range is discussed in Section 3.1. One of these events is shown in Fig. 1, right, where the LDF and energy calibration have been performed using the reconstruction used for the bulk of SD-433 events [38].
@@ 3 Energy assignment of photon and proton events
Unlike hadronic-initiated air showers [54], a data-driven energy calibration for photons is impractical because no photons have yet been detected above a few $\sim 10^{15}$ eV. Furthermore, given the depleted muon content in photon-initiated showers, a data-driven energy calibration would overestimate the photon energy because the SD is sensitive to muons. Stated otherwise, a photon and a hadronic primary of the same energy do not generate the same average signal in the WCDs. Consequently, simulated air showers with a sufficiently large probability to generate an event in the SD array are selected in Section 3.1 before being employed to derive a dedicated energy calibration for photon-initiated events [^2] The bias introduced by the muon deficit in hadronic models is not expected to significantly affect an energy calibration based on photon-initiated events given the minimal muonic content in such showers., as described in Section 3.2.
The discrimination method is designed using protons as the only hadronic species, because proton-initiated air showers have a muon content most similar to photon-initiated showers among the hadronic primaries present in the cosmic-ray flux. The energy of simulated proton events is underestimated when employing a data-driven energy calibration, since current high-energy hadronic interaction models appear to underestimate the number of muons produced in air showers [42, 55]. Therefore, a Monte Carlo-based energy assignment is developed in Section 3.3. This approach provides the most conservative background estimation, as discussed in Section 4.
Because the nature of the primary cosmic rays is unknown in the data, one cannot apply separate energy calibrations for hadron- and photon-initiated showers. A unified energy scale is thus essential for accurately comparing events initiated by different primary species. To tackle this problem, the photon-equivalent energy scale is developed in Section 3.4.
@@ 3.1 Trigger efficiency of photon and proton primaries
Figure 2 depicts the trigger efficiency for the simulated photon and proton events. Air showers initiated by hadronic primaries, such as protons, reach the ground with a significantly more prominent muonic component than those initiated by photon primaries due to the decay of charged pions and kaons produced during hadronic interactions in the atmosphere. This increased muonic content is crucial for trigger generation, especially as the zenith angle increases and atmospheric attenuation of the electromagnetic particles becomes more significant. The trigger efficiency for proton-initiated air showers is thus consistently higher than that for photon events, with the difference becoming more pronounced at larger zenith angles.
@@ 3.2 Photon-initiated events
As described in Section 2, the primary energy is calibrated using the signal interpolated at 250 m from the shower axis, $S(250)$. This value is determined by evaluating a parametrized LDF fitted to best match the observed signals in each triggered SD station. To achieve an accurate estimate of $S(250)$, the slope parameter of the LDF is parametrized as detailed in Appendix A. We utilize only one of the ten realizations of simulated events produced for each air shower to account for shower-to-shower fluctuations in the energy calibration. A power-law model is used to describe the relationship between $S(250)$ and $E_{\text{MC}}$ [63]:
@@
The power-law index, $\alpha$, is found for the subset of events in each of eight zenith-angle bins employing a $\chi^{2}$ minimization. As shown in Fig. 3, left, $\alpha$ is mildly dependent on the zenith angle with a relationship that can be modeled as:
@@
The three free parameters are listed in Table 1. The power-law index $\alpha$ is slightly below unity for vertical events due to the higher likelihood of these showers arriving at the observation level before reaching their maximum development, especially at the highest energies, compared to non-vertical events. The ratio $S(250)/E^{\alpha}$ decreases with the zenith angle, as shown in Fig. 3, right, reflecting the atmospheric attenuation of electromagnetic air showers. Based on the universality of the electromagnetic longitudinal development, the signal-to-energy ratio can be described using a Gaisser-Hillas function [64]:
@@
Figure 3: Left: The power-law index from Eq. 3.1 as a function of zenith angle. The solid line represents the model of Eq. 3.2. Right: The average ratio between shower size and primary energy raised to the power of $\alpha$, in terms of the zenith angle. The fitted model of Eq. 3.3 is represented as the solid line.
The free parameters, listed in Table 1, are estimated via a $\chi^{2}$ minimization to the average ratio. Particularly, the parameter $g_{0}=(38.9\pm 0.1)\,\mathrm{VEM}$ represents the expected signal measured by the SD-433 at 250 m from the axis of a shower initiated by a photon primary of $10^{17}$ eV with a zenith angle of 25 .
The calibration performance is evaluated by examining the relative difference between the reconstructed energy, $E_{\text{rec}}$, and the simulated energy, $E_{\text{MC}}$. This method reconstructs the photon energy with a bias of less than $2\%$, as shown in Fig. 4, top. The resolution of the calibration is defined as the standard deviation of the distribution of the relative difference between $E_{\text{rec}}$ and $E_{\text{MC}}$. As shown in Fig. 4, bottom, the zenith-integrated resolution remains nearly constant at around $12\%$.
Figure 4: The bias and resolution of the energy calibration for photon events defined by Eq. 3.1.
@@
Table 1: The parameters modeling the attenuation of the shower size and the energy calibration, defined by Eqs. 3.2 and 3.3.
@@ 3.3 Proton-initiated events
in analogy with Eq. 3.1. The attenuation function model for proton events, inspired by recent analyses with the SD-433 data [37, 38], is
@@
We select a reference zenith angle of 30 , as it corresponds to the median zenith angle of the simulated proton events, ensuring a representative selection. A combined fit of the attenuation function, $g^{\prime}(\theta)$, and the power-law index, $\alpha^{\prime}$, to the $S(250)$ is conducted. In this process, the likelihood of observing the samples of $S(250)$ is maximized by tuning the mentioned free parameters assuming an underlying Gaussian probability density function describing the theoretical shower size. The standard deviation of the predicted shower size is taken as the uncertainty coming from the event reconstruction. The estimated values for the parameters involved in Eqs. 3.4 and 3.5 are summarized in Table 2.
@@
Table 2: The parameters in Eqs. 3.4 and 3.5 estimated through the maximum likelihood method on the $S(250)$ for proton-initiated events.
The bias and resolution provide insight into the reliability and precision of the calibration. As shown in the top panel of Fig. 5, the reconstructed energy exhibits a bias below $2\%$ for energies above $\sim 10^{16.7}$ eV, while the resolution follows the expected trend with the primary energy, improving from $18\%$ to $13\%$ across the energy range under study, as presented in the bottom panel of Fig. 5.
Figure 5: The bias (top) and resolution (bottom) of the energy calibration for proton events defined by Eq. 3.4.
@@ 3.4 The photon-equivalent energy scale
In the case of simulated proton events, the energy assignment tailored for photon primaries described in Section 3.2 leads to a systematic overestimation of the reconstructed energy by $\sim 15\%$, as shown by the black markers in Fig. 6, left. This is due to the muonic component, larger in hadronic showers than in electromagnetic ones, in particular at 250 m from the shower axis. Furthermore, because proton and photon showers attenuate differently in the atmosphere, one can see in the figure a clear dispersion in the reconstructed energy across angular bins. To account for the different attenuation of hadron- and photon-initiated air showers, we define the photon-equivalent energy scale, $E_{\gamma,\text{eq}}$. This scale is constructed using the attenuation curve derived from simulated proton events, i.e., from Eq. 3.5, and the energy power-law index obtained for photon events, i.e., from Eq. 3.1. Thus a reconstructed event with a shower size $S(250)$ can be assigned an energy $E_{\gamma,\text{eq}}$ independently of the primary species and for both simulations and data.
The application of this energy scale to simulated proton events results in a bias of $10\%$ at $10^{16.7}\,\mathrm{eV}$ decreasing to $5\%$ at the highest energies, as shown in Fig. 6, right. Importantly, this bias shows no significant angular dependence, making $E_{\gamma,\text{eq}}$ a robust scale across different zenith angles. On the other hand, photon-initiated events, when analyzed using this scale, are assigned underestimated energies. The bias changes from $-10\%$ to $-15\%$ in the energy range of interest, as displayed in Fig. 6, right. Consequently, in a given $E_{\gamma,\text{eq}}$ bin, simulated photon events are mixed with proton events that have true lower energies, and thus lower muon content. This mixing leads to a conservative estimation of the separation between hadron- and photon-initiated events based on the muon content that scales with the primary energy. Therefore, the discrimination method described in the next section is tailored and optimized in such a conservative scenario.
Figure 6: The relative difference between the reconstructed and simulated energy in terms of the latter. Left: Events are initiated by proton primaries, where the energy is assigned using the procedure described in Section 3.2. Right: Events are initiated by photon and proton primaries, where the energy is assigned using the photon-equivalent energy scale (see text for details).

### v3

@@ 1 Introduction
In this study, we extend the search for primary photons done at the Auger Observatory towards lower energies, reaching down to $5\times 10^{16}$ eV. The extension is made possible by a dense array of $19$ WCDs deployed over 2 km2 and the direct measurements of the air-shower muonic component by an array of buried scintillators deployed in the same region, as discussed in Section˜2. The Monte Carlo-driven reconstruction of photon-initiated events and the definition of a unique energy scale to treat both photon and hadronic events are presented in Section˜3. The photon-hadron discrimination is performed using an observable, $M_{b}$, defined as the event-wise weighted sum of the muon density measurements. The power of this photon-hadron discriminator is presented in-depth in Section˜4. The selection criteria considered in data are described in Section˜5. The results of the photon search are presented in Section˜6.
@@ 2 Detection systems and data
In the configuration used in this work, eleven stations of the SD-433 array are co-located with stations of the Underground Muon Detector (UMD), as illustrated in Fig.˜1, left. Each UMD station consists of three modules, each containing $64$ plastic scintillator bars measuring 400 cm in length, 4 cm in width and 1 cm in thickness providing a sensitive area of $\sim$10 m2 [42, 43]. The central station on the array’s western edge contains six modules, providing a sensitive area of $\sim$50 m2, compared to the standard 30 m2 area in all other stations. All modules are buried under 2.3 m of soil equivalent to a vertical shielding of 540 g/cm2, that filters out most of the particles produced in the air except muons with kinetic energies $\gtrsim$1 GeV. When a charged particle passes through the module, the scintillation material emits photons that are collected by a 1.2 mm wavelength-shifter optical fiber and conducted to an array of $64$ silicon photomultipliers (SiPMs) Hamamatsu S13081-050CS [44]. Front-end electronics convert the analog pulses of each SiPM into binary signals, effectively translating the current pulses into a sequence of digital boolean samples. These output signals are sampled by a Field-Programmable Gate Array at 320 MHz, corresponding to a sampling interval of 3.125 ns, resulting in a binary trace of $2048$ bits stored in the front-end memory. The back-end electronics handles all calibration, control, and monitoring tasks [45]. Additionally, the surface electronics, common to all UMD station modules, interfaces with the SD electronics to check for a trigger and retrieve and transfer the binary traces upon an event data request.
Figure 1: Left: A schematic view of the SD-433 array. Solid black positions are equipped with buried scintillators. Events acquired by the three highlighted hexagons are considered for the photon search. Right: One of the events acquired on February 27, 2021, part of the selected data set. The reconstructed energy is $(4.1\pm 0.1)\times 10^{17}$ eV, obtained with a dedicated data energy scale [38], with a zenith angle of $(26.3\pm 0.3)$. Dashed lines indicate the fitted lateral distribution functions (see Section˜3).
@@
To reconstruct the incoming direction of the primary particle and the lateral distribution function (LDF) [48] of the air showers we use the signals from the WCDs. The LDF is described by the Nishimura-Kamata-Greisen (NKG) function [49, 50] from which we extract the shower size, $S(r_{\text{ref}})$, defined as the value at an optimal distance, $r_{\text{ref}}$. For an array spacing of 433 m, $r_{\text{ref}}$ is chosen as 250 m to minimize the uncertainties of the signal due to the imperfect knowledge of the functional form of the LDF in individual events [51]. The shower size, $S(250)$, corrected for the attenuation in the atmosphere serves as an observable that can be calibrated against the primary energy. Although for most events detected with SD-433 (hadronic showers) a data-driven energy calibration is employed [38, 52], the energy calibration for photon events is instead performed using detailed Monte Carlo simulations, as discussed in Section˜3.
@@
The photon search strategy designed for this work relies on the SD-433 data to measure the air showers, providing information on the primary energy and shower geometry, and on the direct measurement of the air-shower muon content, as recorded by the UMD stations. To ensure a sufficient sampling of the air-shower muon content, leading to a suitable discrimination power as discussed in Section˜5, we select events acquired by the three highlighted hexagons of Fig.˜1, left. The data for this analysis were recorded from December 17, 2020, when the central UMD stations in the three employed hexagons were commissioned, to March 31, 2022, before the upgrade of the $19$ stations with new electronics [53].
@@
The flux of events acquired by the SD hexagons is expected to remain constant, apart from a negligible seasonal modulation. As the number of detected events follows Poisson statistics, identifying unstable periods in the data acquisition of the SD-433 involves searching for time intervals between consecutive events incompatible with the Poisson expectation [^1] For example, consecutive events separated more than 1.1 hours when the seven hexagons of the array are operative are not compatible with a Poisson probability at a $99\%$ confidence level and hence identified as an unstable period.. The same methodology is employed for the UMD modules, based on the rate of 6T5 events acquired by the SD-433 during which a UMD module is registered as active. Unstable periods for each UMD module are identified by analyzing the time intervals between consecutive appearances in the data. After subtracting unstable periods for both detection systems based on the arrival time of 6T5 events, the final data set is composed of 5T5 and 6T5 events recorded during ${\sim}15.5\,\mathrm{months}$, equivalent to $8\,\mathrm{months}$ of data acquisition with the three hexagons operating under ideal conditions. The final data set is composed of 15,919 events with energies above $10^{16.7}$ eV and zenith angles up to $52\,\mathrm{{}^{\circ}}$. The energy threshold is derived with the Monte Carlo study presented in Section˜3.4, while the zenith angle range is discussed in Section˜3.1. One of these events is shown in Fig.˜1, right, where the LDF and energy calibration have been performed using the reconstruction used for the bulk of SD-433 events [38].
@@ 3 Energy assignment of photon and proton events
Unlike hadronic-initiated air showers [54], a data-driven energy calibration for photons is impractical because no photons have yet been detected above a few $\sim 10^{15}$ eV. Furthermore, given the depleted muon content in photon-initiated showers, a data-driven energy calibration would overestimate the photon energy because the SD is sensitive to muons. Stated otherwise, a photon and a hadronic primary of the same energy do not generate the same average signal in the WCDs. Consequently, simulated air showers with a sufficiently large probability to generate an event in the SD array are selected in Section˜3.1 before being employed to derive a dedicated energy calibration for photon-initiated events [^2] The bias introduced by the muon deficit in hadronic models is not expected to significantly affect an energy calibration based on photon-initiated events given the minimal muonic content in such showers., as described in Section˜3.2.
The discrimination method is designed using protons as the only hadronic species, because proton-initiated air showers have a muon content most similar to photon-initiated showers among the hadronic primaries present in the cosmic-ray flux. The energy of simulated proton events is underestimated when employing a data-driven energy calibration, since current high-energy hadronic interaction models appear to underestimate the number of muons produced in air showers [42, 55]. Therefore, a Monte Carlo-based energy assignment is developed in Section˜3.3. This approach provides the most conservative background estimation, as discussed in Section˜4.
Because the nature of the primary cosmic rays is unknown in the data, one cannot apply separate energy calibrations for hadron- and photon-initiated showers. A unified energy scale is thus essential for accurately comparing events initiated by different primary species. To tackle this problem, the photon-equivalent energy scale is developed in Section˜3.4.
@@ 3.1 Trigger efficiency of photon and proton primaries
Figure˜2 depicts the trigger efficiency for the simulated photon and proton events. Air showers initiated by hadronic primaries, such as protons, reach the ground with a significantly more prominent muonic component than those initiated by photon primaries due to the decay of charged pions and kaons produced during hadronic interactions in the atmosphere. This increased muonic content is crucial for trigger generation, especially as the zenith angle increases and atmospheric attenuation of the electromagnetic particles becomes more significant. The trigger efficiency for proton-initiated air showers is thus consistently higher than that for photon events, with the difference becoming more pronounced at larger zenith angles.
@@ 3.2 Photon-initiated events
As described in Section˜2, the primary energy is calibrated using the signal interpolated at 250 m from the shower axis, $S(250)$. This value is determined by evaluating a parametrized LDF fitted to best match the observed signals in each triggered SD station. To achieve an accurate estimate of $S(250)$, the slope parameter of the LDF is parametrized as detailed in Appendix˜A. We utilize only one of the ten realizations of simulated events produced for each air shower to account for shower-to-shower fluctuations in the energy calibration. A power-law model is used to describe the relationship between $S(250)$ and $E_{\text{MC}}$ [63]:
@@
The power-law index, $\alpha$, is found for the subset of events in each of eight zenith-angle bins employing a $\chi^{2}$ minimization. As shown in Fig.˜3, left, $\alpha$ is mildly dependent on the zenith angle with a relationship that can be modeled as:
@@
The three free parameters are listed in Table˜1. The power-law index $\alpha$ is slightly below unity for vertical events due to the higher likelihood of these showers arriving at the observation level before reaching their maximum development, especially at the highest energies, compared to non-vertical events. The ratio $S(250)/E^{\alpha}$ decreases with the zenith angle, as shown in Fig.˜3, right, reflecting the atmospheric attenuation of electromagnetic air showers. Based on the universality of the electromagnetic longitudinal development, the signal-to-energy ratio can be described using a Gaisser-Hillas function [64]:
@@
Figure 3: Left: The power-law index from Eq.˜3.1 as a function of zenith angle. The solid line represents the model of Eq.˜3.2. Right: The average ratio between shower size and primary energy raised to the power of $\alpha$, in terms of the zenith angle. The fitted model of Eq.˜3.3 is represented as the solid line.
The free parameters, listed in Table˜1, are estimated via a $\chi^{2}$ minimization to the average ratio. Particularly, the parameter $g_{0}=(38.9\pm 0.1)\,\mathrm{VEM}$ represents the expected signal measured by the SD-433 at 250 m from the axis of a shower initiated by a photon primary of $10^{17}$ eV with a zenith angle of 25 .
The calibration performance is evaluated by examining the relative difference between the reconstructed energy, $E_{\text{rec}}$, and the simulated energy, $E_{\text{MC}}$. This method reconstructs the photon energy with a bias of less than $2\%$, as shown in Fig.˜4, top. The resolution of the calibration is defined as the standard deviation of the distribution of the relative difference between $E_{\text{rec}}$ and $E_{\text{MC}}$. As shown in Fig.˜4, bottom, the zenith-integrated resolution remains nearly constant at around $12\%$.
Figure 4: The bias and resolution of the energy calibration for photon events defined by Eq.˜3.1.
@@
Table 1: The parameters modeling the attenuation of the shower size and the energy calibration, defined by Eqs.˜3.2 and $3.$
@@ 3.3 Proton-initiated events
in analogy with Eq.˜3.1. The attenuation function model for proton events, inspired by recent analyses with the SD-433 data [37, 38], is
@@
We select a reference zenith angle of 30 , as it corresponds to the median zenith angle of the simulated proton events, ensuring a representative selection. A combined fit of the attenuation function, $g^{\prime}(\theta)$, and the power-law index, $\alpha^{\prime}$, to the $S(250)$ is conducted. In this process, the likelihood of observing the samples of $S(250)$ is maximized by tuning the mentioned free parameters assuming an underlying Gaussian probability density function describing the theoretical shower size. The standard deviation of the predicted shower size is taken as the uncertainty coming from the event reconstruction. The estimated values for the parameters involved in Eqs.˜3.4 and 3.5 are summarized in Table˜2.
@@
Table 2: The parameters in Eqs.˜3.4 and 3.5 estimated through the maximum likelihood method on the $S(250)$ for proton-initiated events.
The bias and resolution provide insight into the reliability and precision of the calibration. As shown in the top panel of Fig.˜5, the reconstructed energy exhibits a bias below $2\%$ for energies above $\sim 10^{16.7}$ eV, while the resolution follows the expected trend with the primary energy, improving from $18\%$ to $13\%$ across the energy range under study, as presented in the bottom panel of Fig.˜5.
Figure 5: The bias (top) and resolution (bottom) of the energy calibration for proton events defined by Eq.˜3.4.
@@ 3.4 The photon-equivalent energy scale
In the case of simulated proton events, the energy assignment tailored for photon primaries described in Section˜3.2 leads to a systematic overestimation of the reconstructed energy by $\sim 15\%$, as shown by the black markers in Fig.˜6, left. This is due to the muonic component, larger in hadronic showers than in electromagnetic ones, in particular at 250 m from the shower axis. Furthermore, because proton and photon showers attenuate differently in the atmosphere, one can see in the figure a clear dispersion in the reconstructed energy across angular bins. To account for the different attenuation of hadron- and photon-initiated air showers, we define the photon-equivalent energy scale, $E_{\gamma,\text{eq}}$. This scale is constructed using the attenuation curve derived from simulated proton events, i.e., from Eq.˜3.5, and the energy power-law index obtained for photon events, i.e., from Eq.˜3.1. Thus a reconstructed event with a shower size $S(250)$ can be assigned an energy $E_{\gamma,\text{eq}}$ independently of the primary species and for both simulations and data.
The application of this energy scale to simulated proton events results in a bias of $10\%$ at $10^{16.7}\,\mathrm{eV}$ decreasing to $5\%$ at the highest energies, as shown in Fig.˜6, right. Importantly, this bias shows no significant angular dependence, making $E_{\gamma,\text{eq}}$ a robust scale across different zenith angles. On the other hand, photon-initiated events, when analyzed using this scale, are assigned underestimated energies. The bias changes from $10\%$ to $15\%$ in the energy range of interest, as displayed in Fig.˜6, right. Consequently, in a given $E_{\gamma,\text{eq}}$ bin, simulated photon events are mixed with proton events that have true lower energies, and thus lower muon content. This mixing leads to a conservative estimation of the separation between hadron- and photon-initiated events based on the muon content that scales with the primary energy. Therefore, the discrimination method described in the next section is tailored and optimized in such a conservative scenario.
Figure 6: The relative difference between the reconstructed and simulated energy in terms of the latter. Left: Events are initiated by proton primaries, where the energy is assigned using the procedure described in Section˜3.2. Right: Events are initiated by photon and proton primaries, where the energy is assigned using the photon-equivalent energy scale (see text for details).